# DRS-OSS: LLM-Driven Diff Risk Scoring Tool for PR Risk Prediction

### v1

@@ Abstract.
DRS-OSS integrates directly into developer workflows through: (i) a FastAPI gateway and LLM microservices for scalable inference, (ii) a React-based dashboard for manual diff analysis, and (iii) a GitHub App that posts risk labels and confidence scores on pull requests. The system delivers real-time, reproducible risk feedback, and is released with a full replication package including fine-tuning scripts, deployment artifacts (ReplicationPackage), and source code (GitRepo). We additionally provide the tool’s website (ToolWebsite), along with a demonstration video showcasing the end-to-end workflow (DemoVideo).
@@ 1. Introduction
Modern release pipelines increasingly demand not only fast delivery but also automated protection against defects and outages. Diff Risk Scoring (DRS) operationalizes this by asking, for each outgoing diff (pull request), how likely is this change to cause a defect if shipped? For example, at Meta, DRS models assign a calibrated risk score to every diff and use that score to adjust which diffs are merged or gated: low-risk diffs pass with minimal friction, while the riskiest diffs (e.g., top $5\%$) may have a short code freeze or receive additional reviews (Abreu2024).
Conceptually, DRS is a practical instantiation of just-in-time defect prediction (JIT-DP) in the release-deployment setting. JIT-DP reframes defect prediction at the change granularity so that risky commits can be surfaced while the author’s context is still fresh, reducing effort compared to coarser file/package models (mockus2000jitdp; Kamei2013). Classic JIT-DP formulations treat this as a binary classification problem over commits or diffs, producing a risk score and/or label for each change that can then be consumed by downstream tools, such as gating policies.
Beyond flagging risk, JIT-DP outputs can drive several software-engineering actions: (i) review prioritization: route scarce reviewer attention to the riskiest changes or assign expert reviewers first (Huang2018EffortAware); (ii) test selection/prioritization: trigger a focused subset of tests for the change, reducing CI cost while preserving failure detection (Machalica2018PTS; Srikanth2016RiskPrioritization); and (iii) policy-based gating: apply risk-aligned merge gates (e.g., flag the top-$X\%$ riskiest diffs) to reduce severe production incidents (Abreu2024).
In this work, we present DRS-OSS, a DRS tool built for open source projects. Our system is built around a fine-tuned Llama 3.1 8B sequence classifier trained on the ApacheJIT dataset (Keshavarz2022ApacheJIT) of defect-inducing and clean commits. The model consumes long-context representations that combine commit messages, structured diffs, and change metrics as special tokens, and is trained using parameter-efficient adaptation, 4-bit QLoRA, and DeepSpeed ZeRO-3 CPU offloading, enabling 22k-token contexts while consuming only 20 GBs on GPU VRAM. The resulting risk scores can then be converted into practical gating policies for code freeze.
@@
We provide a complete replication package containing all model fine-tuning scripts (ReplicationPackage) and our full implementation of the DRS-OSS service and GitHub integration (GitRepo).
@@ 2. Model Training
This section describes how we train the Llama-based binary classifier that powers DRS-OSS. We choose Llama as our base model because it is pre-trained jointly on natural language and source code, making it well-suited to reason over commit messages and diffs, and because Llama-family models have already been shown to outperform classical baselines for diff risk scoring in production at Meta (Abreu2024).
We fine-tune the pretrained Llama 3.1 8B for binary sequence classification by attaching a dedicated classifier head in place of the original language-modeling head (llama-seq-cls-hf). Concretely, we remove the LM token-distribution layer and introduce a trainable linear projection that maps a single hidden-state vector to a scalar logit. During the forward pass, the transformer produces a hidden state for every token; we take the hidden state of the final non-pad token as a pooled representation of the entire diff, reflecting the model’s contextualized understanding of the sequence. This vector is fed through the linear projection to obtain a raw score, which is then converted to a probability via a sigmoid.
We use ApacheJIT (Keshavarz2022ApacheJIT) as our dataset with an 80/10/10 chronological split (train/validation/test) to avoid temporal leakage, ensuring that the model is trained only on past commits and never exposed to commits from the future.
The inputs to the model are structured with special delimiters to separate commit message, code diff, and change metrics (size, entropy, files touched, etc. (Keshavarz2022ApacheJIT)). We add these delimiters as new tokens to the model during fine-tuning so that it can learn their structural role These tokens are shown in Fig. 2. Since LLMs cannot inherently understand and compare numerical values, we feed change metrics as special tokens to the model. The special tokens are added to the tokenizer vocabulary and the embedding matrix is correspondingly resized. Placing metrics at the beginning gives the encoder early access to coarse change signals while still allowing the model to attend across message and diff content. After normalization, bucketing, and structuring, the metrics section of each sample has the following format:
@@
Class imbalance is addressed with undersampling of the majority class using a 0.7 sampling ratio during batch construction. This reduces gradient dominance from abundant negatives while keeping sufficient diversity of “clean” changes for robust thresholding later (Huang2018EffortAware).
Long diffs demand aggressive efficiency measures. We adopt 4-bit weight quantization (Dettmers2023qlora) for training to fit long-context batches on the GPU. To fine-tune efficiently, we use LoRA adapters on attention and MLP modules while keeping the frozen base weights in low precision. Sequences are truncated to a maximum length of 22,000 tokens, which covers 97.1% of commits in our corpus.
Optimization follows a memory-frugal, scalable recipe. We run DeepSpeed ZeRO-3 (Rajbhandari2021deepspeed) with full parameter and optimizer offloading to CPU, enabling large effective batch sizes despite limited GPU VRAM memory In practice, this configuration consumes roughly 100 GB of host RAM during training. Gradients are accumulated across micro-batches to keep step-level statistics stable while using a per-device batch size of 1. Mixed-precision kernels are enabled where available. The model is trained for 3 epochs; on our hardware budget (a 20 GB NVIDIA A100, single device in our setup), the end-to-end run completes in roughly two weeks, including periodic validation and checkpointing. Overall, this setup allows us to fine-tune DRS-OSS with minimal computational resources while still achieving state-of-the-art results on ApacheJIT.
@@ 3. Evaluation
We evaluate and test DRS-OSS on the ApacheJIT benchmark using an 80/10/10 chronological split to avoid temporal leakage (see Section 2). We compare our model against the best JIT-DP baselines which have used ApacheJIT benchmark: JITGNN (keshavarz2024jitgnn), NeuroJIT (lee2024NeuroJIT), and CodeReviewer (AbuTalib2024).
@@ 4. Results
Prior works (yang2016-effort; fu2017-effort) have used Recall@Top-$k\%$ for $k\!\in\!\{5,10,30\}$, i.e., the fraction of truly risky changes captured if only the riskiest $k\%$ of diffs are quarantined. Table 2 reports Recall@Top-$k\%$ for DRS-OSS. The results imply that gating only the riskiest 30% of diffs would capture 86.4% of defect-inducing commits with minimal developer disruption.
@@ 5.1. Model and API
Figure 2. Commit structuring changes the unified diff into a simpler format that can be better understood by LLMs. It also reduces the number of input tokens in the sequence.:
We provide two inference modes, both implemented using Hugging Face pipeline API (wolf-etal-2020-transformers):
@@ 5.2. Web UI
We illustrate Web UI usage on a real change from apache/hive that fixes an out-of-memory error. Bug-fix changes, especially those touching resource management and error-handling paths, are known to be error-prone due to their complexity and cross-cutting effects (Yin2011fixbug). Fig. 3 highlights how JIT scoring is displayed in web UI.
Figure 3. Demo result on the Hive OOM-fix commit: derived label (risky/safe) and confidence for the predicted label.:
@@ 5.3. GitHub Bot
Figure 4. GitHub bot output on a simple PR with three commits with label and confidence for each commit in the PR.:
@@ 6. Conclusion and Community Feedback
Finally, the tool has attracted external interest: engineering teams at Duolingo, Microsoft, and Uber have reached out regarding replicating or adapting our risk prediction pipeline for their own large-scale continuous integration environments. This feedback highlights the practical value of open, reproducible tooling for diff risk scoring in modern software development workflows.
Acknowledgements.
Portions of the text were revised with assistance from ChatGPT (OpenAI). The authors verified all content for accuracy and take full responsibility for the final manuscript.
@@ A.2. DRS-OSS REST API
The DRS-OSS backend exposes a small REST API hosted at: http://worldofcode.org/drs-api. All endpoints speak HTTP/JSON and are designed to be called either by the DRS-OSS Gateway, the Web UI, or external clients (e.g., scripts, CI jobs).
@@ A.2.1. GET /health
GET /health curl -sS http://worldofcode.org/drs-api/health
@@ A.2.2. POST /seq-cls/predict
POST /seq-cls/predict curl -sS -X POST http://worldofcode.org/drs-api/seq-cls/predict \ -H "Content-Type: application/json" \ -d ’{"diff":"diff --git ...", "commit_message":"Fix NPE"}’
@@ A.2.3. POST /seq-cls/predict_batch
POST /seq-cls/predict_batch curl -sS -X POST http://worldofcode.org/drs-api/seq-cls/predict_batch \ -H "Content-Type: application/json" \ -d ’[{"diff":"...","commit_message":"..."}, {"diff":"...","commit_message":"..."}]’
@@ A.2.4. POST /seq-cls/predict_by_sha
POST /seq-cls/predict_by_sha curl -sS -X POST http://worldofcode.org/drs-api/seq-cls/predict_by_sha \ -H "Content-Type: application/json" \ -d ’{"owner_repo":"facebook/react", "commit_sha":"abc123..."}’
@@ A.2.5. POST /clm/predict
POST /clm/predict curl -sS -X POST http://worldofcode.org/drs-api/clm/predict \ -H "Content-Type: application/json" \ -d ’{"diff":"...","commit_message":"..."}’
@@ A.2.6. POST /clm/predict_by_sha
POST /clm/predict_by_sha curl -sS -X POST http://worldofcode.org/drs-api/clm/predict_by_sha \ -H "Content-Type: application/json" \ -d ’{"owner_repo":"facebook/react", "commit_sha":"abc123..."}’

### v2

# DRS-OSS: Practical Diff Risk Scoring with LLMs

@@ Abstract.
DRS-OSS integrates directly into developer workflows through: (i) a FastAPI gateway and LLM microservices for scalable inference, (ii) a React-based dashboard for manual diff analysis, and (iii) a GitHub App that posts risk labels and confidence scores on pull requests. The system delivers real-time, reproducible risk feedback, and is released with a full replication package including fine-tuning scripts, deployment artifacts (Salehi et al., 2025c), and source code (Salehi et al., 2025a). We additionally provide the tool’s website (Salehi et al., 2025b), along with a demonstration video showcasing the end-to-end workflow (Salehi, 2025).
@@ 1. Introduction
Open-source software (OSS) is a critical component of modern software supply chains, making its quality and reliability a matter of significant concern. However, most OSS projects lack the resources to develop or deploy advanced, industry-grade tools that help prevent defects and outages.
Diff Risk Scoring (DRS) addresses this gap by estimating, for each outgoing diff (pull request), the likelihood that a change will introduce a defect if shipped. For example, at Meta, DRS models assign a calibrated risk score to every diff and use that score to inform merge and gating decisions: low-risk diffs proceed with minimal friction, while the highest-risk changes (e.g., the top $5\%$) may be subject to short code freezes or additional reviews (Abreu et al., 2025).
Conceptually, DRS is a practical instantiation of just-in-time defect prediction (JIT-DP) in the release-deployment setting. JIT-DP reframes defect prediction at the change granularity so that risky commits can be surfaced while the author’s context is still fresh, reducing effort compared to coarser file/package models (Mockus and Weiss, 2000; Kamei et al., 2013). Classic JIT-DP formulations treat this as a binary classification problem over commits or diffs, producing a risk score and/or label for each change that can then be consumed by downstream tools, such as gating policies.
Beyond flagging risk, JIT-DP outputs can drive several software-engineering actions: (i) review prioritization: route scarce reviewer attention to the riskiest changes or assign expert reviewers first (Huang et al., 2019); (ii) test selection/prioritization: trigger a focused subset of tests for the change, reducing CI cost while preserving failure detection (Machalica et al., 2019; Srikanth et al., 2016); and (iii) policy-based gating: apply risk-aligned merge gates (e.g., flag the top-$X\%$ riskiest diffs) to reduce severe production incidents (Abreu et al., 2025).
In this work, we present DRS-OSS, a DRS tool built for open source projects. Our system is built around a fine-tuned Llama 3.1 8B sequence classifier trained on the ApacheJIT dataset (Keshavarz and Nagappan, 2022) of defect-inducing and clean commits. The model consumes long-context representations that combine commit messages, structured diffs, and change metrics as special tokens, and is trained using parameter-efficient adaptation, 4-bit QLoRA, and DeepSpeed ZeRO-3 CPU offloading, enabling 22k-token contexts while consuming only 20 GBs on GPU VRAM. The resulting risk scores can then be converted into practical gating policies for code freeze.
@@
We note additional external interest in the tool from industry practitioners: engineering teams at Duolingo, Microsoft, and Uber have reached out regarding replicating or adapting our risk prediction pipeline for their own large-scale continuous integration environments. This feedback highlights the practical value of open, reproducible tooling for diff risk scoring in modern software development workflows.
We provide a complete replication package containing all model fine-tuning scripts (Salehi et al., 2025c) and our full implementation of the DRS-OSS service and GitHub integration (Salehi et al., 2025a).
@@
Figure 1. End-to-end DRS-OSS architecture.
Diagram showing the GitHub App sending PR events to the Gateway/API, which calls the Llama classifier and returns a risk score rendered in the web UI.
Figure 2. Commit structuring changes the unified diff into a simpler format that can be better understood by LLMs. It also reduces the number of input tokens in the sequence.
An XML text that shows tags for COMMIT_MESSAGE, FILE, REMOVED, and ADDED enclosing various commit and diffs lines.
@@ 2. Model Training
This section describes how we train the Llama-based binary classifier that powers DRS-OSS. We choose Llama as our base model because it is pre-trained jointly on natural language and source code, making it well-suited to reason over commit messages and diffs, and because Llama-family models have already been shown to outperform classical baselines for diff risk scoring in production at Meta (Abreu et al., 2025).
We fine-tune the pretrained Llama 3.1 8B for binary sequence classification by attaching a dedicated classifier head in place of the original language-modeling head (10). Concretely, we remove the LM token-distribution layer and introduce a trainable linear projection that maps a single hidden-state vector to a scalar logit. During the forward pass, the transformer produces a hidden state for every token; we take the hidden state of the final non-pad token as a pooled representation of the entire diff, reflecting the model’s contextualized understanding of the sequence. This vector is fed through the linear projection to obtain a raw score, which is then converted to a probability via a sigmoid.
We use ApacheJIT (Keshavarz and Nagappan, 2022) as our dataset with an 80/10/10 chronological split (train/validation/test) to avoid temporal leakage, ensuring that the model is trained only on past commits and never exposed to commits from the future.
The inputs to the model are structured with special delimiters to separate commit message, code diff, and change metrics (size, entropy, files touched, etc. (Keshavarz and Nagappan, 2022)). We add these delimiters as new tokens to the model during fine-tuning so that it can learn their structural role These tokens are shown in Fig. 2. Since LLMs cannot inherently understand and compare numerical values, we feed change metrics as special tokens to the model. The special tokens are added to the tokenizer vocabulary and the embedding matrix is correspondingly resized. Placing metrics at the beginning gives the encoder early access to coarse change signals while still allowing the model to attend across message and diff content. After normalization, bucketing, and structuring, the metrics section of each sample has the following format:
@@
Class imbalance is addressed with undersampling of the majority class using a 0.7 sampling ratio during batch construction. This reduces gradient dominance from abundant negatives while keeping sufficient diversity of “clean” changes for robust thresholding later (Huang et al., 2019).
Long diffs demand aggressive efficiency measures. We adopt 4-bit weight quantization (Dettmers et al., 2023) for training to fit long-context batches on the GPU. To fine-tune efficiently, we use LoRA adapters on attention and MLP modules while keeping the frozen base weights in low precision. Sequences are truncated to a maximum length of 22,000 tokens, which covers 97.1% of commits in our corpus.
Optimization follows a memory-frugal, scalable recipe. We run DeepSpeed ZeRO-3 (Rajbhandari et al., 2021) with full parameter and optimizer offloading to CPU, enabling large effective batch sizes despite limited GPU VRAM memory In practice, this configuration consumes roughly 100 GB of host RAM during training. Gradients are accumulated across micro-batches to keep step-level statistics stable while using a per-device batch size of 1. Mixed-precision kernels are enabled where available. The model is trained for 3 epochs; on our hardware budget (a 20 GB NVIDIA A100, single device in our setup), the end-to-end run completes in roughly two weeks, including periodic validation and checkpointing. Overall, this setup allows us to fine-tune DRS-OSS with minimal computational resources while still achieving state-of-the-art results on ApacheJIT.
@@ 3. Evaluation
We evaluate and test DRS-OSS on the ApacheJIT benchmark using an 80/10/10 chronological split to avoid temporal leakage (see Section 2). We compare our model against the best JIT-DP baselines which have used ApacheJIT benchmark: JITGNN (Keshavarz and Rodríguez-Pérez, 2024), NeuroJIT (Lee et al., 2024), and CodeReviewer (Abu Talib et al., 2024).
@@ 4. Results
Prior works (Yang et al., 2016; Fu and Menzies, 2017) have used Recall@Top-$k\%$ for $k\!\in\!\{5,10,30\}$, i.e., the fraction of truly risky changes captured if only the riskiest $k\%$ of diffs are quarantined. Table 2 reports Recall@Top-$k\%$ for DRS-OSS. The results imply that gating only the riskiest 30% of diffs would capture 86.4% of defect-inducing commits with minimal developer disruption.
@@ 5.1. Model and API
We provide two inference modes, both implemented using Hugging Face pipeline API (Wolf et al., 2020):
@@
Figure 3. Demo result on the Hive OOM-fix commit: derived label (risky/safe) and confidence for the predicted label.
A compact card showing the commit metadata and the model’s predicted risk probability p, the threshold tau, the resulting label, and whether the CI gate would allow or quarantine the change.
@@ 5.2. Web UI
We illustrate Web UI usage on a real change from apache/hive that fixes an out-of-memory error. Bug-fix changes, especially those touching resource management and error-handling paths, are known to be error-prone due to their complexity and cross-cutting effects (Yin et al., 2011). Fig. 3 highlights how JIT scoring is displayed in web UI.
@@ 5.3. GitHub Bot
Figure 4. GitHub bot output on a simple PR with three commits with label and confidence for each commit in the PR.
A GitHub PR comment showing a title line, a small table with label and confidence for each commit hash.
@@ 6. Conclusion and Community Feedback
Finally, The tool has attracted interest from both industry practitioners and academic researchers seeking to adopt or adapt the risk prediction pipeline, underscoring the practical and research value of open, reproducible diff risk scoring tools.
@@ A.2. DRS-OSS REST API
The DRS-OSS backend exposes a small REST API hosted at: https://worldofcode.org/drs-api. All endpoints speak HTTP/JSON and are designed to be called either by the DRS-OSS Gateway, the Web UI, or external clients (e.g., scripts, CI jobs).
@@ A.2.1. GET /health
GET /health curl -sS https://worldofcode.org/drs-api/health
@@ A.2.2. POST /seq-cls/predict
POST /seq-cls/predict curl -sS -X POST https://worldofcode.org/drs-api/seq-cls/predict \ -H "Content-Type: application/json" \ -d ’{"diff":"diff --git ...", "commit_message":"Fix NPE"}’
@@ A.2.3. POST /seq-cls/predict_batch
POST /seq-cls/predict_batch curl -sS -X POST https://worldofcode.org/drs-api/seq-cls/predict_batch \ -H "Content-Type: application/json" \ -d ’[{"diff":"...","commit_message":"..."}, {"diff":"...","commit_message":"..."}]’
@@ A.2.4. POST /seq-cls/predict_by_sha
POST /seq-cls/predict_by_sha curl -sS -X POST https://worldofcode.org/drs-api/seq-cls/predict_by_sha \ -H "Content-Type: application/json" \ -d ’{"owner_repo":"facebook/react", "commit_sha":"abc123..."}’
@@ A.2.5. POST /clm/predict
POST /clm/predict curl -sS -X POST https://worldofcode.org/drs-api/clm/predict \ -H "Content-Type: application/json" \ -d ’{"diff":"...","commit_message":"..."}’
@@ A.2.6. POST /clm/predict_by_sha
POST /clm/predict_by_sha curl -sS -X POST https://worldofcode.org/drs-api/clm/predict_by_sha \ -H "Content-Type: application/json" \ -d ’{"owner_repo":"facebook/react", "commit_sha":"abc123..."}’